OpenStudy (anonymous):

6:11. Is in proportion with 10:22 true or false

OpenStudy (xishem):

\[\frac{6}{11}=\frac{10}{22}\]Multiply each side by 22...\[(22)\frac{6}{11}=10 \rightarrow 12 \neq 10\]